# ArtFormer: Controllable Generation of Diverse 3D Articulated Objects

### v2

@@ 2.2 Generating Tree-structured Objects
2.3 Learning Representations with Conditional Inputs
 [^1] This subsection is removed in the final version.
Reed et al. [53] present a neural language model trained from scratch for zero-shot visual recognition, enabling accurate image retrieval using text-based representations. Hubert et al. [20] use autoencoders to extract visual-semantic joint embeddings. State-of-the-art methods [54, 47, 2, 1, 24] use a domain-specific embedding layer to learn implicit representations for multi-modal inputs, employing functions like cross-attention to model the joint distribution. Compared to image and text data, 3D-formatted data is less common, leading to extensive research on reconstructing 3D objects from image or text inputs[3, 2, 62, 11, 17]. Chen et al. [3] utilize learning by association and metric learning techniques to learn representations conditioned on text. Liu et al. [32] reconstruct 3D articulated objects from a single image, leveraging DINOv2 [44] and GPT-4o to extract articulation information.
Figure 2: Training Pipline of Shape Prior Mini encoder $\mathcal{E}_{g}$ compresses the geometry latent code $z$ into $c_{g}$, which is then processed by the embedding vectors of codebooks to form $\hat{c}_{g}$. $\hat{c}_{g}$ is the condition for diffusion decoder $\epsilon$. Each sub-part has a semantic label, such as ‘the lid of cup’ or ‘handle of box’. These labels, encoded by the pre-trained text encoder, pass through mini encoder $\mathcal{E}_{s}$. The resultant vector $c_{s}$ is then passed into the diffusion shape prior directly.
@@ 3.2 Diverse and Controllable Shape Prior
Sampling Diverse Shapes. A particularly desirable capability is to generate parts with diverse geometry features given its semantic information. For example, we would like USB caps of different shapes and styles. To enable our model for such capability, we discretize the space of geometry code $c_{g}=\mathcal{E}_{g}(z)$ to allow for sampling. A geometry condition $c_{g}$ is chunked into 4 segments $(c_{g}^{0},c_{g}^{1},c_{g}^{2},c_{g}^{3})$, which is designed to enlarge the capacity of the latent space from $4N$ to $N^{4}$. Then, these 4 segments are used to retrieve $(\hat{c}_{g}^{0},\hat{c}_{g}^{1},\hat{c}_{g}^{2},\hat{c}_{g}^{3})$ from 4 different codebooks $M_{\color[rgb]{1,0,1}t}\in\mathbb{R}^{N\times D}$ ($t$ is the index of codebook) using Gumbel-Softmax sampling:
@@
where $\mathcal{D}_{t,l}$ is the distance matrix and $m_{l}^{{\color[rgb]{1,0,1}t}}\in\mathbb{R}^{D}$ denotes the $l$-th out of $N$ embedding vector in the codebook $M_{\color[rgb]{1,0,1}t}$. The Gumble-Softmax operation is defined as:
@@
where $g_{1},\cdots,g_{k}$ are samples from $\operatorname{Gumbel}(0,1)$ [22]. The softmax temperature $\tau$ controls the diversity of shape prior, which we do not specifically tune in this work. Since $\operatorname{GS}$ sampling is differentiable, the model can still be trained end-to-end.
The diffusion model helps sample from the full continuous space of $z$, given $\hat{c}_{g}$ which is either quantised from $c_{g}$ (produced by $\mathcal{E}_{g}$, during shape prior training) or sample by logits $P_{i}\in\mathbb{R}^{4\times N}$ from codebook $M$ directly (after shape prior training). The shape prior ensure the sampled $z$ is well in distribution. Meanwhile, the stochasticity introduced by discrete sampling improves the generation diversity.
@@ 3.3 Articulation Transformer
After the articulated object is parameterized into a tree structure, each node is treated as a token in the classical transformer architecture. $i$-th token is composed of $[\text{fa}_{i},b_{i},z_{i},j_{i},l_{i}]$, where $\text{fa}_{i}$ is the parent index of $i$-th node. In addition to the index of the parent node $\text{fa}_{i}$, the remaining attributes are concatenated and subsequently processed through an MLP Embedding Mapper. An overview of the articulation transformer is illustrated in Fig. 3.
Tree Position Embedding. In order for the transformer to recognize the specific position of each token, we proposed a novel position encoding scheme specifically designed for tree structures building upon the works of [56, 49, 25]. We first calculate the absolute position encoding $a_{i}{\color[rgb]{1,0,1}\in\mathbb{R}^{64}}$ for each $i$-th node following [49]:
@@
where $\mathcal{A}_{i}=[b_{i},z_{i},j_{i},l_{i}]$, $K$ is the depth of $i$-th node and $\{p_{k}\}_{k=1}^{K}$ is the sequence of index alone the path on the tree from the root to $i$-th node. It push the tokens on the path from the root to a bi-directional GRU [49, 25] to compress the information. We define the position embedding of the $i$-th node $p_{i}{\color[rgb]{1,0,1}\in\mathbb{R}^{1024}}$ to represent the relative position following [56]:
@@
In each iteration, we input all previously generated nodes and predict a child for all of the input nodes, starting from a special start token $\mathcal{S}$, which conditions the generation of the root node. If no child nodes can be added to a current node, a special terminal token $\mathcal{T}$ is outputted. The self-attention layer ensures that the same child token is not repeatedly generated for any node across different prediction iteration. This process can be better understood through the illustration in Fig. 4.
Figure 4: Each blue card represents a round in the predicting process. On each blue card, the left side shows the input given to the model and the expected output. The right side displays the tree structure of the articulated object formed after this predicting round, with green nodes indicating the nodes generated in this round. Orange nodes are terminal nodes.
The decoding procedure terminates when all output tokens are $\mathcal{T}$. We then project the generated tokens back to the format described in Sec. 3.1, i.e., the kinematic characteristics and relations $(b_{i},j_{i},l_{i})$, and conditions $(P,c_{s})_{i}$ for SDF decoding. To obtain the final object, cast joint transforms to rigid transforms to the input of SDF $\Omega$ and extract mesh using the Marching Cube [35] algorithm.
Training Objective. The output of the articulation transformer is a tuples $(o,b,j,l,c_{s},P)_{i}$. Binary cross-entropy loss $L_{o}$ is employed to supervise $o$, which is the logits indicating whether this token is a terminal token. And MSE loss $L_{a}$ is used to supervise the attributes $b,j,l,c_{s}$.
@@ 4.2 Baselines
    Ours-PR: Building on our original model. We perform part retrieval after the iterative decoding as CAGE do for fair comparison.
@@
    Ours-NAPSP: Building on our original model, but the geometry is generated through the shape prior of NAP.
@@ 4.4 Results
Generation Quality and Diversity. We evaluate the generation quality of baselines, as shown in Tab. 1. Since CAGE cannot directly generate geometry, comparisons are divided into two groups. The first includes NAP-128, NAP-768, and our model, which generate geometry features directly. The second approach consists of CAGE and ours-PR, which retrieve the suitable shape from the dataset to generate objects. Our model outperforms NAP on all metrics in the first group, producing more realistic articulated objects with less part interpenetration. In the second group, while CAGE achieves better MMD and POR, indicating superior object-level reconstruction, our model excels in COV and 1-NNA, capturing the overall distribution and generating more diverse objects. HS results in Tab. 1 suggest that our model produces greater diversity and better aligns with text instructions from an ordinary user’s perspective.
@@ 5 Conclusion
We propose a novel method for modeling and generating 3D articulated objects, addressing limitations in diversity and usability. Representing articulated objects as a tree structure with nodes for rigid parts simplifies articulation parameterization, enabling part-level definition and generation. To ensure well-aligned yet diverse outputs, we develop a controllable shape prior and the an Articulation Transformer, which captures articulation features effectively. A tree position embedding layer enhances part relationship modeling, supporting autoregressive generation. Our method achieves state-of-the-art performance, generating high-quality, diverse objects from text or image conditions.
Limitations and Future Work. (1) The limited dataset restricts the application to a small range of object types with few sub-part, preventing the full potential of the approach from being realized. Our model is expected to show more advantage when the number of parts is large ($>10$) and varies (which are rare in existing dataset), due to transformer’s capability to handle long sequences of varying lengths. Future work may explore this capability in large scale. (2) Multi-modal instructions beyond text and images have not yet been explored, such as point cloud or joint structure of expected articulated object. Investigating diverse instruction formats could greatly enhance flexibility and usability of our method in application. (3) Capturing joint quantitative details in the text condition, such as rotation angles, is more challenging than joint type and geometry condition. Further research is needed to improve representation and learning of this data. (4) It is observed that increasing the number of object categories during training leads to a decline in geometry reconstruction quality, likely due to the limited generalization of SDF models. Methods from GenSDF [5] may enhance generalization, and we also recommend using modern 3D representations like 3D Gaussian Splatting [26] in future work.
Acknowledgements. This paper is a pure student work, as all authors are either undergraduate or PhD students. The computational resources were provided by Ningxia Westcloud Computing Technology Co., Ltd. (Damodel), which offered us certain discounts and waivers. Partial financial support was provided by Xiamen University Malaysia.

### v3

@@ 2.2 Generating Tree-structured Objects
Figure 2: Training Pipeline of Shape Prior Mini encoder $\mathcal{E}_{g}$ compresses the geometry latent code $z$ into $c_{g}$, which is then processed by the embedding vectors of codebooks to form $\hat{c}_{g}$. $\hat{c}_{g}$ is the condition for diffusion decoder $\epsilon$. Each sub-part has a semantic label, such as ‘the lid of cup’ or ‘handle of box’. These labels, encoded by the pre-trained text encoder, pass through mini encoder $\mathcal{E}_{s}$. The resultant vector $c_{s}$ is then passed into the diffusion shape prior directly.
@@ 3.2 Diverse and Controllable Shape Prior
Sampling Diverse Shapes. A particularly desirable capability is to generate parts with diverse geometry features given the semantic information. For example, we would like USB caps of different shapes and styles. To enable our model for such capability, we discretize the space of geometry code $c_{g}=\mathcal{E}_{g}(z)$ to allow for sampling. A geometry condition $c_{g}$ is chunked into 4 segments $(c_{g}^{0},c_{g}^{1},c_{g}^{2},c_{g}^{3})$, which is designed to enlarge the capacity of the latent space from $4N$ to $N^{4}$. Then, these 4 segments are used to retrieve $(\hat{c}_{g}^{0},\hat{c}_{g}^{1},\hat{c}_{g}^{2},\hat{c}_{g}^{3})$ from 4 different codebooks $M_{\color[rgb]{1,0,1}t}\in\mathbb{R}^{N\times D}$ ($t$ is the index of codebook) using Gumbel-Softmax sampling:
@@
where $\mathcal{D}_{t,l}$ is the distance matrix and $m_{l}^{{\color[rgb]{1,0,1}t}}\in\mathbb{R}^{D}$ denotes the $l$-th out of $N$ embedding vector in the codebook $M_{\color[rgb]{1,0,1}t}$. The Gumbel-Softmax operation is defined as:
@@
where $g_{1},\cdots,g_{k}$ are samples from ${Gumbel}(0,1)$ [16]. The softmax temperature $\tau$ controls the diversity of shape prior, which we do not specifically tune in this work. Since $\operatorname{GS}$ sampling is differentiable, the model can still be trained end-to-end.
The diffusion model helps sample from the full continuous space of $z$, given $\hat{c}_{g}$ which is either quantised from $c_{g}$ (produced by $\mathcal{E}_{g}$, during shape prior training) or sample by logits $P_{i}\in\mathbb{R}^{4\times N}$ from codebook $M$ directly (after shape prior training). By leveraging the shape prior, the sampled $z$ is guaranteed to align well with the target distribution. Meanwhile, the stochasticity introduced by discrete sampling improves the generation diversity.
@@ 3.3 Articulation Transformer
After the articulated object is parameterized into a tree structure, each node is treated as a token in the classical transformer architecture. The $i$-th token is composed of $[fa_{i},b_{i},z_{i},j_{i},l_{i}]$, where ${fa}_{i}$ is the parent index of the $i$-th node. In addition to the index of the parent node $fa_{i}$, the remaining attributes are concatenated and subsequently processed through an MLP Embedding Mapper. An overview of the articulation transformer is illustrated in Fig. 3.
Tree Position Embedding. In order for the transformer to recognize the specific position of each token, we proposed a novel position encoding scheme specifically designed for tree structures building upon the works of [45, 40, 18]. We first calculate the absolute position encoding $a_{i}{\color[rgb]{1,0,1}\in\mathbb{R}^{64}}$ for each $i$-th node:
@@
where $\mathcal{A}_{i}=[b_{i},z_{i},j_{i},l_{i}]$, $K$ is the depth of $i$-th node and $\{p_{k}\}_{k=1}^{K}$ is the sequence of indices along the path on the tree from the root to $i$-th node. It pushes the tokens on the path from the root to a bi-directional GRU [40, 18] to compress the information. We define the position embedding of the $i$-th node $p_{i}{\color[rgb]{1,0,1}\in\mathbb{R}^{1024}}$ to represent the relative position:
@@
In each iteration, we input all previously generated nodes and predict a child for all of the input nodes, starting from a special start token $\mathcal{S}$, which conditions the generation of the root node. If no child nodes can be added to a current node, a special terminal token $\mathcal{T}$ is outputted. The self-attention layer ensures that the same child token is not repeatedly generated for any node across different prediction iterations. This process can be better understood through the illustration in Fig. 4.
Figure 4: Each blue card represents a round in the predicting process. On each blue card, the left side shows the input given to the model and the expected output. The right side displays the tree structure of the articulated object formed after this prediction round, with green nodes indicating the nodes generated in this round. Orange nodes are terminal nodes.
The decoding procedure terminates when all output tokens are $\mathcal{T}$. We then project the generated tokens back to the format described in Sec. 3.1, i.e., the kinematic characteristics and relations $(b_{i},j_{i},l_{i})$, and conditions $(P,c_{s})_{i}$ for SDF decoding. To obtain the final object, cast joint transforms to rigid transforms to the input of SDF $\Omega$ and extract mesh using the Marching Cubes [28] algorithm.
Training Objective. The output of the articulation transformer are some tuples ($[o,b,j,l,c_{s},P]_{i}$ for the $i$-th node). Binary cross-entropy loss $L_{o}$ is employed to supervise $o$, which is the logits indicating whether this token is a terminal token. MSE loss $L_{a}$ is used to supervise the attributes $b,j,l,c_{s}$.
@@ 4.2 Baselines
    Ours-PR: Building on our original model. We perform part retrieval after iterative decoding, as CAGE does, for a fair comparison.
@@
    Ours-NAPSP: Building on our original model, where the geometry is generated by the shape prior of NAP.
@@ 4.4 Results
Generation Quality and Diversity. We evaluate the generation quality of baselines, as shown in Tab. 1. Since CAGE cannot directly generate geometry, comparisons are divided into two groups. The first includes NAP-128, NAP-768, and our model, which generate geometry features directly. The second approach consists of CAGE and ours-PR, which retrieve the suitable shape from the dataset to generate objects. Our model outperforms NAP on all metrics in the first group, producing more realistic articulated objects with less part interpenetration. In the second group, while CAGE achieves better MMD and POR, indicating superior object-level reconstruction, our model excels in COV and 1-NNA, capturing the overall distribution and generating more diverse objects. HS results in Tab. 1 suggest that our model produces greater diversity and aligns better with text instructions from an ordinary user’s perspective.
@@ 5 Conclusion
We propose a novel method for modeling and generating 3D articulated objects, addressing limitations in diversity and usability. Representing articulated objects as a tree structure with nodes for rigid parts simplifies articulation parameterization, enabling part-level definition and generation. To ensure well-aligned yet diverse outputs, we develop a controllable shape prior and the Articulation Transformer, which captures articulation features effectively. A tree position embedding layer enhances part relationship modeling, supporting autoregressive generation. Our method achieves state-of-the-art performance, generating high-quality, diverse objects from text or image conditions.
Limitations and Future Work. (1) The limited dataset restricts the application to a small range of object types with few sub-part, preventing the full potential of the approach from being realized. Our model is expected to show more advantage when the number of parts is large ($>10$) and varies (which are rare in existing dataset), due to transformer’s capability to handle long sequences of varying lengths. Future work may explore this capability on a large scale. (2) Multi-modal instructions beyond text and images have not yet been explored, such as point cloud or joint structure of expected articulated object. Investigating diverse instruction formats could greatly enhance flexibility and usability of our method in practical application. (3) Capturing joint quantitative details in the text condition, such as rotation angles, is more challenging than joint type and geometry condition. Further research is needed to improve representation and learning of this data. (4) It is observed that increasing the number of object categories during training leads to a decline in geometry reconstruction quality, likely due to the limited generalization of SDF models. Methods from GenSDF [2] may enhance generalization, and we also recommend using modern 3D representations like 3D Gaussian Splatting [19] in future work.
Acknowledgements. This paper is a pure student work, as all authors are either undergraduate or PhD students. Special thanks to Ningxia Westcloud Computing Technology Co., Ltd. for DaModel’s cloud services and student academic discounts, vital to this research. Partial financial support was provided by Xiamen University Malaysia.